how to find the proper divisors of a given number

Grade:11

2 Answers

Sanjeev Malik IIT BHU
12 years ago

let a no be 21600

now 21600=52*25*33

so the powers are 2,5,3

so no of divisors= (2+1)*(5+1)*(3=1)=72 divisors

(  see(2+1): it is actually no of ways of selecting two 5s to get a no=total three ways-->selecting no 5, one 5 and both(all)5  )

( all combinations are made to get a no. which is a divisor0

more u can clearify when you will read permuttions and combinations..

 

 

please approve by clciking "YES" below

Jaya IITK
12 years ago

By Lcm